solve
5x^2=2x-7x^2
3x^2+6x=2x^2-9

Answer by stanbon(75887)

solve
5x^2=2x-7x^2
3x^2+6x=2x^2-9
---
Simplify:
12x^2-2x = 0
2x(6x-1) = 0
x = 0 or x = 1/6
---------------------
x^2 + 6x + 9 = 0
Factor:
(x+3)^2 = 0
x = -3 with multiplicity two
==============================
Cheers,
Stan H.